# AI Gender Bias, Disparities, and Fairness: Does Training Data Matter?

### 17 Dec 2023 v1

@@ 3.5. Equalized Odds
These three methods - Paired t-test, MSG, and Equalized Odds - provide a comprehensive approach to evaluating gender bias, disparities, and fairness in AI-based scoring systems, ensuring the ethicaal uses of AI models. The methodology encompassed a detailed experimental setup with rigorous training and testing phases, coupled with statistical analyses to assess gender issues in AI-based scoring systems comprehensively.
@@ 4.2. Mean Score Gap
To verify our assumption, we designed two experiments to compare the difference of MSGs between human- and machine-graded scores. We tested the three models: Mixed-, male-, and female-trained models, and compared each of them with human-generated MSG.
@@ 4.3.1. BERT Model Equalized Odds
Interpretation: The mixed trained models for both BERT and GPT-3.5 have lower EO values, suggesting more equitable predictions and higher fairness compared to gender-specific models. This outcome supports the hypothesis that training with balanced dataset can reduce the EO gradient, contributing to the development of fair AI models.
@@ 5. Discussion
This study demonstrates that gender unbalanced training datasets doesn’t have traces of gender bias for AI-based scoring systems. The prevalent notion that AI inherently perpetuates gender biases (Bolukbasi et al. (Bolukbasi et al., 2016); Lu et al. (Lu et al., 2020)) is challenged by the outcomes of this study. However, our research shows that models trained on data lacking gender balance might lead to algorithms exacerbating gender disparities. The reduced MSG and lower EO gradients in models trained on mixed datasets highlight the potential of balanced training in addressing gender fairness. This finding is crucial in the context of Hardt et al. (Hardt et al., 2016), who emphasize the importance of equality of opportunity in supervised learning.
In education, where AI’s fairness is paramount (Li et al. (Li et al., 2022)), our study suggests that balanced mixed training datasets have the potential to enhance the equity of AI-based assessments. This approach aligns with the ethical frameworks proposed by Slimi (Slimi, 2023) for AI in higher education, emphasizing the need for inclusive and equitable AI practices.
The concept of pseudo AI bias (Zhai & Krajcik, 2022(Zhai and Krajcik, 2022)) implies that biases in AI are not always overt but can be subtle and deeply ingrained. The evidence of gender neutral behavior of the balanced mixed training model in this study provides a roadmap for addressing such false claims regarding biases in AI. By ensuring diverse representation in training data, AI models can be steered away from inadvertently reinforcing gender stereotypes.
Our findings resonate with the literature advocating for debiasing AI systems (Nadeem et al., 2022(Nadeem et al., 2022); Manresa-Yee & Ramis, 2021(Manresa-Yee and Ramis, 2021)). This study contributes to the broader efforts in creating fair and unbiased AI models by demonstrating that balanced training has more equitable AI outcomes. The approach counters the notion that AI systems are destined to replicate societal biases, offering a practical solution to this pervasive issue.
The results of this study present a compelling argument for the gender issue of balanced mixed training in AI for automatic scoring. This approach addresses the immediate concerns of fairness and equity in AI-based scoring systems and provides evidence against mechanical pseudo-bias in broader AI applications.
@@ Acknowledgment
This study secondary analyzed data from project supported by the Institute of Education Sciences (grant number R305A160219). The authors acknowledge the funding agencies and the project teams for making the data available for analysis. Project is partially supported by NSF (grants numbers 2101104). The findings, conclusions, or opinions herein represent the views of the authors and do not necessarily represent the views of personnel affiliated with the funding agencies.

### 26 Dec 2023 v2

@@ 3.5. Equalized Odds
These three methods - Paired t-test, MSG, and Equalized Odds - provide a comprehensive approach to evaluating gender bias, disparities, and fairness in AI-based scoring systems, ensuring the ethical uses of AI models. The methodology encompassed a detailed experimental setup with rigorous training and testing phases and statistical analyses to assess gender issues in AI-based scoring systems comprehensively.
@@ 4.2. Mean Score Gap
To verify our assumption, we designed two experiments to compare the difference of MSGs between human- and machine-graded scores. We tested the three models, Mixed-, male-, and female-trained models, and compared each of them with human-generated MSG.
@@ 4.3.1. BERT Model Equalized Odds
Interpretation: The mixed-trained models for both BERT and GPT-3.5 have lower EO values, suggesting more equitable predictions and higher fairness compared to gender-specific models. This outcome supports the hypothesis that training with a balanced dataset can reduce the EO gradient, contributing to the development of fair AI models.
@@ 5. Discussion
The research shows that gender bias is absent from gender-unbalanced training datasets for AI-based scoring systems. The results provided evidence to refute the widely held belief that AI inevitably reinforces gender biases (Bolukbasi et al., 2016; Lu et al., 2020). Our findings, however, suggest that algorithms that are trained on data that is imbalanced in terms of gender could exacerbate existing gaps in gender. The promise of balanced training in addressing gender equity is demonstrated by the decreased EO gradients and reduced MSG in models trained on mixed datasets. This discovery is vital in light of Hardt et al. (2016), who highlight the significance of opportunity equality in supervised learning.
Our research indicates that balanced mixed training datasets have the potential to improve the equality of AI-based assessments in education, where the fairness of AI is crucial (Li et al., 2022). This strategy emphasizes the need for inclusive and equitable AI activities, which aligns with the ethical principles Slimi (2023) suggests for AI in higher education.
According to Zhai and Krajcik (2022) concept of pseudo AI bias, biases in AI might be deeply embedded and subtle rather than always obvious. This study’s demonstration of the balanced mixed training model’s gender-neutral behavior offers a path forward for resolving such unfounded allegations about AI biases. Gender stereotypes can be prevented from unintentionally being reinforced by AI models by guaranteeing varied representation in training data.
The literature supporting debiasing AI systems (Manresa-Yee and Ramis, 2021; Nadeem et al., 2022) aligns with our findings. This work shows that balanced training produces more equitable AI outcomes, which advances the larger attempts to develop impartial and fair AI models. The method offers a workable answer to this widespread problem by refuting that AI systems are bound to reproduce societal prejudices.
The study’s findings strongly support the gender issue of balanced mixed training in artificial intelligence for automated scoring. In addition to addressing the pressing issues of equality and fairness in AI-based scoring systems, our method offers proof against mechanical pseudo-bias in more general AI applications.
@@ Acknowledgment
This study secondary analyzed data from a project supported by the Institute of Education Sciences (grant number R305A160219). The authors acknowledge the funding agencies and the project teams for making the data available for analysis. The project is partially supported by NSF (grant number 2101104). The findings, conclusions, or opinions herein represent the authors’ views and do not necessarily represent the views of personnel affiliated with the funding agencies.